# Configurations, Tessellations and Tone Networks

### v4

@@ X Tone Networks for Tetrachords
—D. Tymoczko (2011), A Geometry of Music
We promised that we would return to the construction of tonnetze for tetrachords. The work that has been hitherto carried out in this area offers many insights into the relation of the families of tetrachords to one another but at the same time forces one to realise the daunting nature of the task. In Tristan und Isolde, the unordered minor sixth (or half-diminished seventh) chord (i.e., the Tristan chord or one of its permutations) resolves in no less than eight distinct ways into dominant seventh chords, according to Tymoczko (2012) at p. 302. Transposing the initial chord (as he does) to $G_{\sharp}m^{6}$, the eight listed resolutions are to $E^{7}$, $F^{7}$, $G^{7}$, ${A_{\flat}}^{\!7}$, ${B_{\flat}}^{\!7}$, $B^{7}$, ${D_{\flat}}^{\!7}$, and ${D}^{7}$. This list includes Chopin’s resolution to ${B_{\flat}}^{\!7}$ in the $G$-minor Ballade, discussed earlier, but not the Tchaikovsky resolution to ${E_{\flat}}^{\!7}$ in the Adagio Lamentoso, which nonetheless works quite well as a resolution, and perhaps ought to be added to the list on an honorary basis. On that argument it is tempting to believe that the empirical evidence suggests that any resolution to a dominant seventh is admissible.
Can one do better than this? Our approach will be to adopt a variant of the strategy that we used to create the Archimedean tonnetze. Thus, we consider all transitions from minor sixth chords to dominant sevenths with the property that exactly two tones change. The goal is then to derive a Levi graph for the resulting relations. The problem that immediately arises is that there are too many of such transitions, some of them even admitting tetracycles, such as $\langle{A_{\flat}m}^{6},$ $B^{7},$ ${Dm}^{6},$ $F^{7},$ ${A_{\flat}m}^{6}\rangle$, hence blocking the possibility of a Levi graph. Thus, we need some principle that can be applied uniformly across the class of chords being considered to limit the range of transitions sufficiently to ensure that we get a Levi graph. Our criterion for success will be that the resulting Levi graph is associated with an interesting configuration.
@@
Therefore, let us impose the condition that the dissonant note is unchanged. That implies that two members of the underlying consonant triad have to shift, and we are back to a situation similar to what we considered earlier in connection with the Archimedean tonnetz. The result is constrained by the requirement that a minor sixth must transform into a dominant seventh and vice versa. It is easy to see that there are exactly three transitions that send $Cm^{6}$ to a dominant seventh chord in such a way that the $A$ is retained, namely $Cm^{6}\to A^{7}$,  $Cm^{6}\to B^{7}$, and $Cm^{6}\to D^{7}$. Similarly, there are exactly three ways in which $C^{7}$ can be sent to a minor sixth such that the $B_{\flat}$ is retained, namely $C^{7}\to B_{\flat}m^{6}$,  $C^{7}\to E_{\flat}m^{6}$, and $C^{7}\to C_{\sharp}m^{6}$. Similar transformations exist by transposition for the other minor sixths and dominant sevenths. Putting these together, we obtain a connected bipartite graph of degree three and girth six for the minor sixths and dominant sevenths, shown in Figure 20. Thus, we have a Levi graph and a combinatorial configuration $\{12_{3}\}$.
@@
By counting cycles, one easily determines that this graph is different from the Levi graph of the Eulerian tonnetz. Now, the Eulerian tonnetz has a total of sixteen hexacycles – that is, the four ‘hexatonic’ 3pp-hexacycles and the twelve ‘pitch-retaining’ 2pp-hexacycles. But the Tristan graph admits twenty-four hexacycles – specifically, we have twelve 1pp-hexacycles (hexabeanies) of the form $\langle C^{7},E_{\flat}m^{6},F^{7},G_{\sharp}m^{6},A_{\sharp}^{7},C_{\sharp}m^{6},C^{7}\rangle$ and twelve 2pp-hexacycles (bow ties) of the form $\langle C^{7},E_{\flat}m^{6},F^{7},F_{\sharp}m^{6},D_{\sharp}^{\,7},C_{\sharp}m^{6},C^{7}\rangle$.
@@
Note that we have retained the ‘$p$’ notation for edges that cross the graph, where in the present context one can take it to mean ‘preserve’ (as in ‘preserving the dissonant note’) rather than ‘parallel’. Since our new Levi graph has a different number of hexacycles than that of the Eulerian tonnetz, this means that the two are not isomorphic. On the other hand, since it is known that all the combinatorial $\{12_{3}\}$ are realizable as configurations in the Euclidean plane, one can ask which of the 229 it turns out to be, if it is not the D220. We have been able to identify it as the D228 of Daublebsky von Sterneck (1895). This rather attractive and highly symmetric configuration is plotted in Figure 21. Since they are isomorphic, we can refer to either of the Levi graph in Figure 20 and the configuration shown in Figure 21 as the ‘Tristan tonnetz’.   [^21] Using a metaphysical analogy, one can think of these three abstractions – tonnetz, Levi graph and configuration – as forming a Trinity: with the Configuration as Father, the Graph as Son, and the Tonnetz as Geist. Thus, Father, Son and Geist are distinct and yet inseparable.
@@
A tonnetz for the Tristan genus of dominant seventh chords and minor sixth chords can be constructed by limiting the maps between these chord sets to two-note transitions that preserve the dissonant note of the chord being mapped. The resulting incidence geometry, which is distinct from that of the Eulerian tonnetz, is the configuration $\{12_{3}\}$ in $\mathbb{R}^{2}$ known as the D228 of Daublebsky von Sterneck. The $12$ dominant sevenths are represented by points and the $12$ minor sixths are represented by lines. The $36$ incidence relations between the $12$ points and $12$ lines determine the edges of the associated Levi graph, where the white vertices are dominant sevenths and the black vertices are minor sixths.
Figure 21: The geometric configuration of the Tristan genus tonnetz is the D228 of Daublebsky von Sterneck, for which the corresponding Levi graph appears in Figure 20. This self-dual configuration is one of the four ‘homogeneous’ cases arising among of the 229 configurations of type $\{12_{3}\}$. The group of the configuration is generated by quarter-turn physical rotations of the overall figure and cyclic permutations of the three mutually inscribed squares among each other. The three squares correspond to the three $4p$-octacycles of the Levi graph.
The symmetry group of this Tristan-genus configuration is generated by quarter-turn physical rotations of the overall figure and cyclic permutations of the three mutually inscribed squares among each other. The three squares, which are on an equal footing, correspond to the three $4p$-octacycles of the Levi graph. The musical significance of the squares of the configuration, or equivalently the octacycles of the Levi graph, is that movements along or within these structures give rise some of the characteristic sonorities of late Wagner operas. We mention several examples of these ‘Wagnerian octacycles’.
First, we return to the prelude to Act I of Tristan und Isolde. The reader will no doubt be aware that three different versions the opening motive shown in Figure 5 appear in succession at the outset of the prelude. The first occurrence is as in Figure 5. In the second occurrence, the entire motive is transposed up by a minor third, apart from the first note, which is transposed up only one tone, so the interval between the first and the second note is a major sixth rather than the minor sixth one hears in the first occurrence of the motive. In the third occurrence, the initial rise is again by an interval of a major sixth, but the resolution of the minor sixth chord to a dominant seventh is different from that of the first two occurrences. In more detail, the three resolutions are as follows:
@@
In an octacycle with the structure $\langle 1,2,3,4,5,6,7,8,1\rangle$, where the odd numbers represent minor sixths and the even numbers represent dominant sevenths, we note that the polar opposite of any chord is necessarily a chord of the same mode. Thus, the polar of chord 1 is chord 5, which is half way around the cycle. But for some purposes we may be interested in the most opposite chords of the opposite mode. We can call these the ‘subpolars’; there are two of them. Thus, the subpolars of chord 1 are chords 4 and 6. With this terminology we are ready to look at relations of the chords in the Tristan prelude. We observe that the polar of $G_{\sharp}m^{6}$ along the octacycle is $Dm^{6}$. The subpolars of $G_{\sharp}m^{6}$ are given by $E^{7}$ and $B^{7}$.
@@
Thus, Wagner’s well known first resolution (i) runs along the octacycle beginning at $G_{\sharp}m^{6}$ and resolving to the subpolar $E^{7}$. A glance at Figure 20 shows that resolution (ii) begins on the same octacycle at $Bm^{6}$ and resolves to its subpolar, which is $G^{7}$. Resolution (iii) again begins on the same octacycle, this time at $Fm^{6}$, but resolves to $B^{7}$, which is not on the same octacycle – but this $B^{7}$ nevertheless stands in a definite relation to $Fm^{6}$: in fact, $B^{7}$ is polar to $Fm^{6}$ along the 2pp-hexacycle (a bow tie) that connects these two chords. In summary, all three minor sixths – $G_{\sharp}m^{6}$, $Bm^{6}$, and $Fm^{6}$ – lie on the same octacycle along with two of the resolvent dominant sevenths, $E^{7}$ and $G^{7}$.
@@
But the final resolution starting at $Fm^{6}$ is deflected along one of the three 2pp-hexacycles containing that chord, namely the one that passes through $E^{7}$ and $Dm^{6}$, to reach its polar $B^{7}$, which we recall is the other subpolar of the original $G_{\sharp}m^{6}$. Thus, all of the chords of this profound opening gesture of the Tristan prelude stand in a definite and rather subtle set of relations to one another on the tonnetz.  [^22] This passage at the beginning of the opera seems to have sprung fully formed from the composer’s mind. One is reminded of the analysis of William Blake’s Infant Sorrow by Roman Jakobson (1970), who after similar consideration of the intricate syntactic and morphological symmetries of Blake’s short poem comes to the conclusion that it couldn’t have been assembled piece by piece.
@@
While we are on the topic of Tristan, let us recall the resolutions used by Tchaikovsky and Chopin that we had given a tentative discussion earlier, in Section IV. How do these resolutions square up on the Tristan tonnetz? In the case of Tchaikovsky, we observe that his resolution (in Wagner’s pitches) from $G_{\sharp}m^{6}$ to $D_{\sharp}^{7}$ at the beginning of the Adagio Lamentoso is polar along the unique 1pp-hexacycle connecting these chords. In the case of Chopin’s resolution of the Tristan chord from $G_{\sharp}m^{6}$ to $B\flat^{7}$ at the transition to the più animato section of the Ballade in $G$-minor, we find interestingly that this $B\flat^{7}$ is adjacent to $G_{\sharp}m^{6}$ on the Tristan tonnetz. So there seems to be a kind of tonnetz logic to Chopin’s choice that eludes classical analysis.  [^23] We have approached these resolutions both from a triadic perspective and now also from a tetrachordal perspective. Why both? We are working at the boundary of art and science. It is of the nature of works of art that one can never quite pin them down. On the one hand, they resist analysis, and on the other hand they will readily embrace multiple modes of analysis. Each such approach reveals yet another aspect of the artwork. If the object under consideration yields completely to any one single mode of analysis, then it isn’t art. In this respect, we seem to be in agreement with Tomoczko (2011) at p. 304: ‘ .  .  .  I am not proposing a simple method or rule for doing musical analysis: there is no royal road to musical understanding, geometrical or otherwise.’
Here we can offer a modest proposal towards a resolution a controversy in the theory of the tonnetz concerning the notion of ‘distance’. The idea that a measure of distance between chords can be formulated by counting the minimum number of steps along the tonnetz that it takes to get from one chord to another has been explored by many authors, but never in a completely satisfactory way.  [^24] Tomoczko (2011) at Appendix C, offers a critique of such tonnetz-based distance measures. Does it make sense to say that $E_{\flat}M$ is closer to $Gm$ (one step away on the Eulerian tonnetz) than is $DM$ (three steps away)? Then why does Bach (as would most of his contemporaries) conclude the first movement of his Fantasy and Fugue in $G$-minor for organ with a cadence from $DM$ to $Gm$, rather than somehow, say, resolving from $E_{\flat}M$?
The answer may lie in the fact that $DM$ is polar to $Gm$ along the unique 2pp-hexacycle of the Eulerian tonnetz, and the idea that the relation of polarity along a cycle is somehow connected with the emotion of ‘closure’. The two significance relations in this respect would appear to be adjacency and polarity.  [^25] We enter not the debate on whether such psychological notions can be meaningfully incorporated into music theory in a systematic way; but there is one positive lesson that we might learn from geometry in this connection. This is the point that although distance is not really part of the story in projective geometry, which is the area of geometry to which the theory of configurations properly belongs, the notions of adjacency, co-cyclicity and polarity are. The proposal then is that if we confine our psychological remarks to those concerning relations of adjacency, co-cyclicity and polarity, then our analysis is more likely to make sense.
@@
In conclusion, we consider another Wagnerian example of the resolution of minor sixths to dominant sevenths, this time in an octacycle of the Tristan tonnetz from Götterdämmerung. In the final section of the opera, at the beginning of the scene where Brünnhilde orders the vassals to pile up logs to create a great pyre around the dead Siegfried’s body that once lit will also consume Brünnhilde along with the whole of Valhalla, there is an extraordinary sequence of chords of the Tristan genus. The sequence is as follows, extending over a total of twelve measures, beginning at the piano $A^{7}$, building up to a climax at the forte $F_{\sharp}M$:
@@
The final transition $C_{\sharp}^{\,7}\to F_{\sharp}M$ can be heard as an authentic cadence designed as a ‘wrap up’. What interests us is the sequence of six chords leading up to this. Looking at the Tristan tonnetz, one sees that all six tetrads of the sequence lie sequentially on the same octacycle, starting at $A^{7}$. Lewin (1996) and Cohn (2012), at pp. 156-157, present analyses of this sequence from their own perspectives, but our view is that the characterization of this important progression of minor sixths and dominant sevenths as belonging to a single octacycle on the Tristan tonnetz of Figures 20 and 21 makes for a very satisfactory account of the matter.
In fact, the next chord along the octacycle is an $F_{\sharp}^{\,7}$ and we can reasonably take the view that the purpose of Wagner’s intervention of the off-cycle $C_{\sharp}^{\,7}$ just before the would-be $F_{\sharp}^{\,7}$ is to put the brakes on the progression, allowing it to halt at the triadic reduced chord $F_{\sharp}M$ rather than continuing with a fully tetradic $F_{\sharp}^{\,7}$, which would have demanded yet another move to a minor sixth.
The geometrical structure of this progression is particularly striking in the path that it makes as it transverses the configuration in Figure 21. Beginning at the point $A^{7}$ in the lower left corner, the progression progresses ever stately across the line $B_{\flat}m^{6}$, where it lingers for a moment, then on to the point $C^{7}$ and then to the line $D_{\flat}m^{6}$ where again it lingers, then on to $E_{\flat}^{\,7}$ and finally to $Em^{6}$, at which point the brakes are applied, and the terminus is reached with $F_{\sharp}M$ substituting for $F_{\sharp}^{\,7}$. Thus we march right around the outer square of the configuration, omitting only the last line – for had that line been included, then we would be taken back to the beginning.
One could even imagine an alternate reality wherein Wagner, in a moment of enthusiasm, decides to repeat the entire sequence to draw out this scene a bit further, dropping the $C_{\sharp}^{\,7}$, playing the full $F_{\sharp}^{\,7}$, then adding in just one more chord – a $Gm^{6}$ – that would take the sequence back to the $A^{7}$ for a repeat, in a kind of perpetuum mobile.
@@
$B_{\flat}m^{6}\to D_{\flat}m^{6}\to(G_{\flat}m^{6})\to Em^{6}\to G^{\circ 7}.$ (19)
@@
Specifically, we see that the initial minor sixth $B_{\flat}m^{6}$ and the final minor sixth $Em^{6}$ are in polar opposition on the Hamiltonian cycle in Figure 20, as they are in the triadic case, with the first resolution to the ‘Tchaikovsky chord’ $D_{\flat}m^{6}$ halfway between the starting point and the ending point, and the second resolution $G_{\flat}m^{6}$ at a ‘staging post’ one black vertex along the cycle towards the end point. The first, second and fourth chords of (19) lie on the octacycle containing the beginning and ending points. The final resolution to $G^{\circ 7}$, which acts again as a wrap up, is only one semitone away from the would-be continuation $Gm^{6}$, which is another quarter turn around the octacycle. We hope to pursue this and other such analyses further elsewhere.

### v5

@@ X Tone Networks for Tetrachords
—D. Tymoczko, A Geometry of Music
We promised that we would return to the construction of tonnetze for tetrachords. The work that has been hitherto carried out in this area offers many insights into the relation of the families of tetrachords to one another but at the same time forces one to realize the daunting nature of the task. In Tristan und Isolde, the unordered minor sixth (or half-diminished seventh) chord (i.e., the Tristan chord or one of its permutations) resolves in no less than eight distinct ways into dominant seventh chords, according to Tymoczko (2012) at p. 302. If we transpose the initial chord (as does Tymoczko) to $G_{\sharp}m^{6}$, the eight listed resolutions are to the following dominant sevenths: $E^{7},F^{7},G^{7},{A_{\flat}}^{\!7},{B_{\flat}}^{\!7},B^{7},{D_{\flat}}^{\!7},{D}^{7}$. This list includes Chopin’s resolution to ${B_{\flat}}^{\!7}$ in the $G$-minor Ballade, discussed earlier, but not Tchaikovsky’s resolution to ${E_{\flat}}^{\!7}$ in the Adagio Lamentoso, which nonetheless seems to work quite well as a resolution. Perhaps it too ought to be added to the list on an honorary basis? On that argument it is tempting to believe that almost any resolution to a dominant seventh is acceptable. Is there any logic to Wagner’s choice of admissible resolutions?
Our approach to the matter of the Tristan resolutions will be to adopt a variant of the strategy that we used to create the Archimedean tonnetze. Thus, we consider all transitions from minor sixth chords to dominant sevenths with the property that exactly two tones change. The goal is then to derive a Levi graph for the resulting relations. The problem that immediately arises is that there are too many of such transitions, some of them even admitting tetracycles, such as $\langle{A_{\flat}m}^{6},$ $B^{7},$ ${Dm}^{6},$ $F^{7},$ ${A_{\flat}m}^{6}\rangle$, hence blocking the possibility of a Levi graph. Thus, we need some principle that can be applied uniformly across the class of chords being considered to limit the range of transitions sufficiently to ensure that we get a Levi graph. Our criterion for success will be that the resulting Levi graph is associated with an interesting configuration.
@@
Figure 22: A Tristan-genus tonnetz. The twelve minor sixth chords and twelve dominant seventh chords can be arranged in a Levi graph with the incidence geometry of a configuration $\{12_{3}\}$. Each minor sixth chord is connected to three dominant seventh chords and each dominant seventh is connected to three minor sixths. At each transformation two notes are retained in the new chord, one being the ‘dissonant’ note of the original tetrachord, which may or may not be a dissonance in the new one. Edges that cross the tonnetz preserve the dissonant note of the tetrad as a dissonance in the new tetrad. For example, the dissonant $F$ of $G_{\sharp}m^{6}$ is mapped to the dissonant $F$ of $G^{7}$.
@@
Therefore, let us impose the condition that the dissonant note is unchanged. That implies that two members of the underlying consonant triad have to shift, and we are back to a situation similar to what we considered earlier in connection with the Archimedean tonnetze. The result is constrained by the requirement that a minor sixth must transform into a dominant seventh and vice versa.
It is easy to see that there are exactly three transitions that send $Cm^{6}$ to a dominant seventh chord in such a way that the $A$ is retained, namely $Cm^{6}\to A^{7}$,  $Cm^{6}\to B^{7}$, and $Cm^{6}\to D^{7}$. Likewise, there are exactly three ways in which $C^{7}$ can be sent to a minor sixth such that the $B_{\flat}$ is retained, namely $C^{7}\to B_{\flat}m^{6}$,  $C^{7}\to E_{\flat}m^{6}$, and $C^{7}\to C_{\sharp}m^{6}$. Similar transformations exist by transposition for the other minor sixths and dominant sevenths. Putting these together, we obtain a connected bipartite graph of degree three and girth six for the minor sixths and dominant sevenths, shown in Figure 22. Remarkably, we thus obtain a Levi graph and a self-dual configuration $\{12_{3}\}$.
@@
By counting cycles, one easily determines that this graph is different from the Levi graph of the Eulerian tonnetz. Now, the Eulerian tonnetz has a total of sixteen hexacycles – that is, the four ‘hexatonic’ 3pp-hexacycles and the twelve ‘pitch-retaining’ 2pp-hexacycles. But the Tristan graph in Figure 22 admits twenty-four hexacycles – specifically, we have twelve 1pp-hexacycles (hexabeanies) of the form $\langle C^{7},E_{\flat}m^{6},F^{7},G_{\sharp}m^{6},A_{\sharp}^{7},C_{\sharp}m^{6},C^{7}\rangle$ and twelve 2pp-hexacycles (bow ties) of the form $\langle C^{7},E_{\flat}m^{6},F^{7},F_{\sharp}m^{6},D_{\sharp}^{\,7},C_{\sharp}m^{6},C^{7}\rangle$.
@@
Note that we have retained the ‘$p$’ notation for edges that cross the graph, where in the present context one can take it to mean ‘preserve’ (as in ‘preserving the dissonant note’) rather than ‘parallel’. Since our new Levi graph has a different number of hexacycles than that of the Eulerian tonnetz, this means that the two are not isomorphic. On the other hand, since it is known that all the combinatorial $\{12_{3}\}$ configurations are realizable in the Euclidean plane, one can ask which of the 229 it turns out to be, if it is not the D222. We have been able to identify it as the D228 of Daublebsky von Sterneck (1895). This rather attractive and highly symmetric configuration is plotted in Figure 23. Since they are isomorphic, we can refer to either of the Levi graph in Figure 22 and the configuration shown in Figure 23 as the ‘Tristan tonnetz’.   [^28] Using a metaphysical analogy, one can think of these three abstractions – tonnetz, Levi graph and configuration – as forming a Trinity: with the Configuration as Father, the Graph as Son, and the Tonnetz as Geist. Thus, Father, Son and Geist are distinct and yet inseparable.
@@
A tonnetz for the Tristan genus of dominant seventh chords and minor sixth chords can be constructed by limiting the maps between these chord sets to two-note transitions that preserve the dissonant note of the chord being mapped. The resulting incidence geometry, which is distinct from that of the Eulerian tonnetz, is the self-dual configuration $\{12_{3}\}$ in $\mathbb{R}^{2}$ known as the D228 of Daublebsky von Sterneck. The $12$ dominant sevenths are represented by points and the $12$ minor sixths are represented by lines. The $36$ incidence relations between the $12$ points and $12$ lines determine the edges of the associated Levi graph, where the white vertices are dominant sevenths and the black vertices are minor sixths.
Figure 23: The geometric configuration of the Tristan-genus tonnetz is the D228 of Daublebsky von Sterneck, for which the corresponding Levi graph appears in Figure 22. This self-dual configuration is one of the four ‘homogeneous’ cases arising among of the 229 configurations of type $\{12_{3}\}$. The group of the configuration is generated by quarter-turn physical rotations of the overall figure and cyclic permutations of the three mutually inscribed squares among each other. The three squares correspond to the three $4p$-octacycles of the Levi graph.
The symmetry group of this Tristan-genus configuration is generated by quarter-turn physical rotations of the overall figure and cyclic permutations of the three mutually inscribed squares among each other. The three squares, which are on an equal footing, correspond to the three $4p$-octacycles of the Levi graph. The musical significance of these three mutually inscribed quadrilaterals of the configuration is that movements within these structures give rise some of the characteristic sonorities of late Wagner operas.
We shall look at several examples of the use of these ‘Wagnerian octacycles’. First, we return to the prelude to Act I of Tristan und Isolde. Three different versions the opening motif shown in Figure 5 appear in succession at the outset of the prelude. The first occurrence is as in Figure 5. In the second occurrence, the entire motif is transposed up by a minor third, apart from the first note, which is transposed up only one tone, so the interval between the first and the second note is a major sixth rather than the minor sixth one hears in the first occurrence of the motif. In the third occurrence, the initial rise is again by an interval of a major sixth, but the resolution of the minor sixth chord to a dominant seventh is different from that of the first two occurrences. In more detail, the three resolutions are as follows:
@@
In an octacycle with the structure $\langle 1,2,3,4,5,6,7,8,1\rangle$, where the odd numbers represent minor sixths and the even numbers represent dominant sevenths, we note that the polar of any chord is necessarily a chord of the same mode. Thus, the polar of chord 1 is chord 5, which is half way around the cycle. But for some purposes we may be interested in the ‘most opposite’ chords of the opposite mode. We can call these the ‘subpolars’; there are two of them. Thus, the subpolars of chord 1 are chords 4 and 6. With this terminology we are ready to look at relations of the chords in the Tristan prelude. In the tables of cycle numbers shown in the Appendix we see that the Tristan tonnetz admits twelve 3pp-octacycles and three 4pp-octacycles. We observe that the polar of $G_{\sharp}m^{6}$ along the unique 4pp-octacycle containing that chord is $Dm^{6}$. The subpolars of $G_{\sharp}m^{6}$ are $E^{7}$ and $B^{7}$.
@@
Thus, Wagner’s well known first resolution (i) runs along the octacycle beginning at $G_{\sharp}m^{6}$ and resolves at the subpolar $E^{7}$. A glance at Figure 22 shows that resolution (ii) begins on the same octacycle at $Bm^{6}$ and resolves to its subpolar, which is $G^{7}$.
Resolution (iii) again begins on the same octacycle, this time at $Fm^{6}$, but resolves to $B^{7}$, which is not on the same octacycle – but this $B^{7}$ nevertheless stands in a definite relation to $Fm^{6}$: in fact, $B^{7}$ is polar to $Fm^{6}$ along the 2pp-hexacycle (a bow tie) that connects these two chords.
@@
In summary, all three minor sixths – $G_{\sharp}m^{6}$, $Bm^{6}$ and $Fm^{6}$ – lie on the same octacycle of the Tristan tonnetz, along with two of the resolvent dominant sevenths, $E^{7}$ and $G^{7}$. But the final resolution starting at $Fm^{6}$ is deflected along one of the three 2pp-hexacycles containing that chord, namely the one that passes through $E^{7}$ and $Dm^{6}$, to reach its polar $B^{7}$, which we recall is the other subpolar of the original $G_{\sharp}m^{6}$. Hence, all of the chords of this profound opening gesture of the Tristan prelude stand in a definite and rather subtle set of relations to one another on the Tristan tonnetz.  [^29] This passage at the beginning of the opera seems to have sprung fully formed from the composer’s mind. One is reminded of the remarkable analysis of William Blake’s Infant Sorrow carried out by the linguist Roman Jakobson (1970), who after similar consideration of the intricate syntactic and morphological symmetries of Blake’s short poem comes to the conclusion that it couldn’t have been assembled piece by piece.
Another example of the use of octacycles can be found in Act 2, Scene 2 of Tristan und Isolde – the ‘Liebesnacht’. The duet is marked by two warnings from Brangäne (‘Habet acht!’) that night will soon give way to day, following which first Tristan and then Isolde plead for death (‘Lass mich sterben!’). This section is accompanied by a well-known sequence of six three-chord progressions, each extending from minor sixth to minor sixth to dominant seventh, visiting all twelve of the minor sixth chords and half of the dominant seventh chords along the way. Let’s see what this looks like on the Tristan tonnetz. The six progressions, in two groups of three, separated in pitch by a semitone, are as follows:
$(i)\hskip 7.11317ptA_{\flat}m^{6}\to Dm^{6}\to C^{7},\hskip 14.22636pt(ii)\hskip 7.11317ptB_{\flat}m^{6}\to Em^{6}\to D^{7},\hskip 14.22636pt(iii)\hskip 7.11317ptCm^{6}\to F_{\sharp}m^{6}\to E^{7},\hskip 14.22636pt$ (12)
$(iv)\hskip 7.11317ptAm^{6}\to E_{\flat}m^{6}\to C_{\sharp}^{7},\hskip 14.22636pt(v)\hskip 7.11317ptBm^{6}\to Fm^{6}\to D_{\sharp}^{7},\hskip 14.22636pt(vi)\hskip 7.11317ptC_{\sharp}m^{6}\to Gm^{6}\to F^{7}.\hskip 14.22636pt$ (13)
In each progression the two minor sixth chords are polar opposites on the Hamiltonian cycle of the Tristan tonnetz (i.e., twelve edges apart around the rim of the tonnetz) and are also polar opposites on the unique 4pp-octacycle containing them (four edges apart). For instance, $G_{\sharp}m^{6}$ and $Dm^{6}$ are polar on the perimeter Hamiltonian cycle and on the 4pp-octacycle
$\langle A_{\flat}m^{6},G^{7},Fm^{6},E^{7},Dm^{6},C_{\sharp}^{\,7},Bm^{6},A_{\sharp}^{\,7},A_{\flat}m^{6}\rangle.$ (14)
Moreover, again in each of the six progressions, the first minor sixth chord and the resulting dominant seventh chord are simultaneously polar opposites (three edges apart) on the $2p$-hexacycle containing them and on the two $1p$-hexacycles containing them.
Next, we return to Tymoczko’s list of the eight dominant seventh chords admitted in the opera Tristan. Can we make sense of this list in the light of what we have learned? First, we observe that the names of these chords spell out the notes of the octatonic scale $O_{23}$. We note that the list splits into two sets, differing by a semitone: namely, $E^{7},G^{7},{B_{\flat}}^{\!7},{D_{\flat}}^{\!7}$, and ${D}^{7},F^{7},{A_{\flat}}^{\!7},B^{7}$. But these are the seventh chords that belong to precisely two of the three 4pp-octacycles of the Tristan tonnetz. So now we understand which chords are missing from the list: they are the four dominant sevenths belonging to the excluded octacycle.
@@
While we are on the topic of Tristan und Isolde, let us recall the resolutions used by Tchaikovsky and Chopin that we had tentatively discussed in Section IV. How do these resolutions square up on the Tristan tonnetz? In the case of Tchaikovsky, we observe that this resolution from $G_{\sharp}m^{6}$ to $D_{\sharp}^{7}$ (in Wagner’s pitches) at the beginning of the Adagio Lamentoso is polar along the unique 1pp-hexacycle connecting these chords. In the case of Chopin’s resolution from $G_{\sharp}m^{6}$ to $B\flat^{7}$ at the transition to the più animato section of the Ballade in $G$-minor, we find that this $B\flat^{7}$ is adjacent to $G_{\sharp}m^{6}$ on the Tristan tonnetz. So indeed there seems to be a kind of tonnetz logic to Chopin’s choice that eludes classical analysis.  [^30] We have approached these resolutions both from a triadic perspective and now from a tetrachordal perspective. Why both? We are working at the boundary of art and science. It is of the nature of works of art that one can never quite pin them down. On the one hand, they resist analysis, and on the other hand they will readily embrace multiple modes of analysis. Each such approach reveals yet another aspect of the artwork. One might even say that if the object under consideration yields completely to any one mode of analysis, then it isn’t art. In this respect, we seem to be in agreement with Tomoczko (2011) at p. 304: ‘ .  .  .  I am not proposing a simple method or rule for doing musical analysis: there is no royal road to musical understanding, geometrical or otherwise.’ We enter not the debate on whether psychological notions can be meaningfully incorporated into music theory in a mathematically systematic way. All such endeavors remain speculative, though in our view certainly worth attempting. Nonetheless, even the idea of the perceived ‘distance’ between chords is fraught with difficulty. Tomoczko (2011) offers a critique of tonnetz-based distance measures at Appendix C.
@@
We turn to consider another Wagnerian example of the resolution of minor sixths to dominant sevenths, this time in an octacycle of the Tristan tonnetz from Götterdämmerung. In the final section of the opera, at the beginning of the scene where Brünnhilde orders the vassals to pile up logs to create a great funeral pyre around the dead Siegfried’s body, there is an extraordinary well-known sequence of chords of the Tristan genus. The sequence is as follows, extending over a total of twelve measures, beginning softly at the $A^{7}$, building up to a climax at the $F_{\sharp}M$:
@@
The final transition $C_{\sharp}^{\,7}\to F_{\sharp}M$ can be heard as an authentic cadence designed as a ‘wrap up’. What interests us is the sequence of six chords leading up to this. Looking at the Tristan tonnetz, one sees that all six tetrads of the sequence lie sequentially on the same octacycle, starting at $A^{7}$. Lewin (1996) and Cohn (2012), at pp. 156-157, present analyses of this sequence from their own perspectives, but our view is that the characterization of this important progression of minor sixths and dominant sevenths as belonging to a single octacycle on the Tristan tonnetz of Figures 22 and 23 makes for a very satisfactory account of the matter. In fact, the next chord along the octacycle is an $F_{\sharp}^{\,7}$ and we can reasonably take the view that the purpose of Wagner’s intervention of the off-cycle $C_{\sharp}^{\,7}$ just before the would-be $F_{\sharp}^{\,7}$ is to put the brakes on the progression, allowing it to halt at the triadic reduced chord $F_{\sharp}M$ rather than continuing with a fully tetradic $F_{\sharp}^{\,7}$, which would have demanded yet another move to a minor sixth.
The geometrical structure of this progression is particularly striking in the path that it makes as it transverses the configuration in Figure 23. Beginning at the point $A^{7}$ in the lower left corner, the progression progresses in an ever stately way along the line $B_{\flat}m^{6}$, where it lingers, then on to the point $C^{7}$ and then to the line $D_{\flat}m^{6}$ where again it lingers, then on to $E_{\flat}^{\,7}$ and finally to $Em^{6}$, at which point the brakes are applied, and the terminus is reached with $F_{\sharp}M$ substituting for $F_{\sharp}^{\,7}$, as we indicated. Thus we march right around the outer perimeter of the configuration, omitting only the last line – indeed, for had that line been included, then we would be taken back to the beginning. One could even imagine a kind of alternate reality wherein Wagner, in a moment of enthusiasm, decides to repeat the entire sequence to draw out this scene a bit further, dropping the $C_{\sharp}^{\,7}$, playing the full $F_{\sharp}^{\,7}$, then adding in just one more chord – a $Gm^{6}$ – that would take the sequence back to the $A^{7}$ for a repeat, in a kind of perpetuum mobile.  [^31] From this perspective, the progression in this final scene of Götterdämmerung can be understood as a musical realization of a certain quadrilateral in the geometric configuration of the Tristan tonnetz. The music forms a ‘representation’ of the geometry. Perhaps that is the way to think about it.
@@
$B_{\flat}m^{6}\to D_{\flat}m^{6}\to G_{\flat}m^{6}\to Em^{6}\to G^{\circ 7}.$ (16)
@@
Specifically, we see that the initial minor sixth $B_{\flat}m^{6}$ and the final minor sixth $Em^{6}$ are in polar opposition on the Hamiltonian cycle in Figure 22, as they are in the triadic case, with the first resolution to the ‘Tchaikovsky chord’ $D_{\flat}m^{6}$ halfway between the starting point and the ending point, and the second resolution $G_{\flat}m^{6}$ at a ‘staging post’ one black vertex along the cycle towards the end point. The first, second and fourth chords of (16) lie on the octacycle containing the beginning and ending points. The final resolution to $G^{\circ 7}$, which acts again as a wrap up, is only one semitone away from the would-be continuation $Gm^{6}$, which is another quarter turn around the octacycle.